OpenStudy (dottiecat):

Which step should be completed first to solve the equation -2=3z+4? A. Add 2 to both sides. B. Subtract 4 from both sides. C. Divide both sides by 3. D. Add 4 to both sides.

OpenStudy (griefer101):

i think that it would be subtract 4 from both sides

OpenStudy (dottiecat):

@baru @ijlal are you two there? ;3

OpenStudy (serenity2001):

B

OpenStudy (dottiecat):

@Qwertty123 @Roxy.girl

OpenStudy (dottiecat):

Hello? :l

OpenStudy (serenity2001):

It's B -2=3z+4 -4 -4 -6 = 3z /3 /3 -2 = z

OpenStudy (dottiecat):

Thanks @serenity2001 :)

OpenStudy (serenity2001):

np

OpenStudy (dottiecat):

Mind helping me with 4 more?

OpenStudy (serenity2001):

Alright

OpenStudy (dottiecat):

kk, lemme make it, and ill tag you! :)